# Anti-Slip AI-Driven Model-Free Control with Global Exponential Stability in Skid-Steering Robots

### v1

@@ Abstract
Index Terms:
Robust control, mobile robots, heavy-duty machinery, neural networks.
@@ I Introduction
The past few years have been marked by significant global changes driven by events such as a worldwide pandemic, supply chain disruptions, resource shortages, and climate change. These factors have accelerated technological shifts, particularly through advancements in robotization and artificial intelligence (AI), which are transforming various business and industrial sectors [1]. One such sector undergoing transformation is the heavy-duty machinery industry, which now requires new insights into autonomy, reliability, and process efficiency for its future development [2]. Robotizing heavy-duty machinery requires advanced solutions that encompass both theoretical developments and application-oriented approaches, along with real-world experimental validations, to enhance automation and safety. As a result, the development of skid-steering heavy-duty robots (SSHDRs) without a separate steering mechanism has gained increasing demand in industries such as construction, agriculture, and mining. These robot platforms with differential drive systems can perform smallest radius turns by moving one set of wheels forward and the other in reverse, allowing them to pivot within their own length. This capability makes them highly maneuverable and ideal for compact, powerful, and agile operations in confined spaces, but also prone to sliding [3][4]. For instance, a four-wheel SSHDR prototype operating on rough terrain is shown in Fig. 1.
@@
These SSHDRs commonly involve a continuous, real-time engagement of different stages within a framework [5], as shown in Fig. 2.
@@
Dynamic controllers are the lowest executive system, responsible for robustly adjusting the multi-stage mechanical actuator mechanism to generate sufficient motion and ensuring stability and goal achievement. The actuation systems integrated into SSHDRs typically consist of multi-component mechanisms designed to transform energy from the power source into mechanical motion. This process involves energy excitation, high-torque motor operation, high-ratio gearing, and, ultimately, the generation of wheel movement. The inherent complexity and nonlinearities of this multi-stage actuation system pose significant challenges in accurately modeling the SSHDR’s dynamics [6]. Consequently, designing a model-based controller is significantly complicated by substantial modeling uncertainties and external disturbances. Hence, traditional model-free controllers remain widely used in the heavy-duty machinery industry due to their simplicity and lower computational demands, despite advancements in intelligent control methodologies. However, the performance of the common model-free controllers in higher-than-second-order systems is often limited, leading to oscillations, poor performance, or instability [7]. In addition, the safety and stability of SSHDRs on uneven terrain are significantly affected by wheel slippage, which arises from factors such as terrain irregularities, imbalanced load distribution, suboptimal wheel design, dynamic forces during movement, and limitations in the robot’s control mechanisms that may hinder timely adjustments [8]. Notably, despite the critical impact of wheel slippage, research on this issue across all types of heavy-duty robots has received limited attention over the past two decades [9][10]. Various terrains with differing slip ratios must be thoroughly analyzed to ensure the stability of the robot’s autonomous controller before integrating it with mapping systems and obstacle-free trajectory generators. Hence, [11] integrated a coordinated adaptive robust control scheme with a torque allocation technique to solve the chattering phenomenon and wheel slip compensation. However, it assumed the model of the studied skid-steering robot is accurately known. To improve the dynamic controller performance and tracking accuracy, [12] developed a traditional controller, the cascaded proportional–integral–derivative (PID) method, while compensating for the sideslip angles. However, most studies have been built on the assumption that all four wheels of the studied robot can be controlled independently, which makes it easier to address the slipping issue and compensate for it.
@@
This paper addresses the problem of robust stabilization for a four-wheel SSHDR system, in which each two-wheel side can be controlled with a single actuator. The system is characterized by a highly nonlinear multi-component actuation model and operates on severely slippery terrains. In contrast to [11] and [12], the proposed control framework is entirely independent of the SSHDR model and interactive multi-stage actuation mechanisms. In addition, unlike [5], the robustness of the proposed control is guaranteed, adaptively ensuring operation on rough terrains with varying slip ratios. Within the proposed control framework, the functional modeling of the wheel dynamics is approximated using radial basis function neural networks (RBFNNs), where, in contrast to [13] and [14], the NN weights, approximation errors, and their norms are not required to be known. The proposed novel adaptive algorithm updates these values and is also responsible for compensating for disturbances caused by wheel slippage. By incorporating the online-updated values of uncertain functions and parameters into the control framework, for the first time in SSHDRs, model-free tracking of the commanded motion with global exponential stability is ensured. Global exponential stability, a strong form of stability in applications, implying uniform, asymptotic stability with exponential convergence rate, holds for all initial conditions. This work lays the foundation for advancing reliable autonomous navigation in heavy-duty machinery undergoing robotization, as shown in Fig. 2.
@@ II Differential Drive System Dynamics
where $i\in\{R,L\}$, $V_{i}$ is the linear velocity of each wheeled side, and $\bm{V}=[V_{R},V_{L}]^{\top}\in\mathbb{R}^{2}\rightarrow\mathbb{R}$. $U_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is the control input, $d_{i}(\cdot):\mathbb{R}\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ represents the uncertain multi-stage actuation model of the robot, $g_{i}\in\mathbb{R}^{+}$ is the unknown control coefficient due to the inertia of the robot, and $\Delta_{i}(t):\mathbb{R}\rightarrow\mathbb{R}$ represents external disturbances. The disturbance structure under slippage can be changed to $\Delta_{i}(t)=-{\mu_{s}}_{i}(s)F_{id}$, where $F_{id}\in\mathbb{R}\rightarrow\mathbb{R}$ represents external disturbance under no-slip conditions, and ${\mu_{s}}_{i}(s)=1+s_{i}$ denotes the functional slippage effect derived from [15]. The slip ratio of the wheel ($s_{i}\in\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$) can be defined as [16]:
@@ III Rbfnn-Based NN-RMFC
$\Phi_{ik}\left(\bm{V}\right)=\exp\left(-\frac{\left\|\bm{V}-{\varepsilon}_{ik}\right\|^{2}}{\alpha_{ik}^{2}}\right),\quad k=1,2,\ldots,L_{i}$ (6)
[…]
@@
Definition 1 [17]. Consider a nonlinear system as follows:
$\dot{x}=f(x,t)\quad x\left(t_{0}\right)=x_{0}\quad x\in\mathbb{R}^{n}$ (9)
@@ IV Stability Proof
[…]
@@ VI-A Simulation Results in Slippery Terrains
Building upon the work of [18], an external, independent method for self-assessment is employed in this study. Specifically, the Simscape MultibodyTM simulation environment is utilized to validate the proposed control scheme and to examine the impact of the introduced slip condition on the final solution. The simulation focuses on a four-wheel SSHDR with a differential drive system (two wheels on each side), where the selection of chassis configurations is flexible, as the governing equations are not constrained to any particular design or set of physical properties. As shown in Table I, derived from [19, 20, 21, 22, 23], during the execution, we considered randomly non-functional values of tires’ slip ratios under different environmental conditions for Eqs. (2) and (1). Although these are approximations, they are sufficient to demonstrate the control strategy’s effectiveness.
@@
Dry asphalt $0.05-0.40$ $0.05-0.20$
Wet asphalt $0.05-0.80$ $0.05-0.50$
Gravel $0.05-0.5$ $0.05-0.40$
Mud $0.05-0.70$ $0.05-0.50$
Ice $0.05-0.90$ $0.05-0.75$
@@
The proposed control system including Eqs. (5)–(8) was applied to the simulated model, considering slip ratios provided in Table I, while the control design parameters were chosen as follows: $\kappa_{i}=1.2$, $\epsilon_{i}=0.04$, $\sigma_{i}=11.5$, $\gamma_{i}=1.6$, $L_{i}=9$, $\varepsilon_{i,k}=[2\hskip 2.84544pt{rand}-1,2\hskip 2.84544pt{rand}-1]^{\top}$, and $\alpha_{i,k}=0.13$, where $rand$ outputs a variable between 00 and $1$. The platform tracking errors for different operational surfaces are shown in Fig. 8, validating the efficacy of the proposed control. This figure illustrates that the control-implemented robot exhibited a higher overshoot and longer settling time on high-slip-ratio surfaces while achieving a slightly lower steady-state error after the settling time. The zoomed-in section of the figure shows that dry asphalt had the fastest settling time, followed by gravel, mud, and wet asphalt, respectively, with the icy surface having the slowest settling time.
@@ VI-B Experimental Results in Extremely Slippery and Snowy Terrains
The control design parameters were chosen as: $\kappa_{i}=1.9$, $\epsilon_{i}=0.08$, $\sigma_{i}=17.1$, $\gamma_{i}=3.6$, $L_{i}=8$, $\varepsilon_{i,k}=[2\hskip 2.84544pt{rand}-1,2\hskip 2.84544pt{rand}-1]^{\top}$, and $\alpha_{i,k}=0.15$, where $rand$ outputs a variable between 00 and $1$. Fig. 11 illustrates the tracking control performance by depicting the actual measured velocity of each side, commanded by the proposed controller, along with the reference velocity generated by a human-assisted joystick. Note that a difference in the sign of the velocity indicates that the robot is pivoting, as the operator demands.
@@
The comparative results in Table II indicate that, overall, the proposed NN-RMFC demonstrates better performance for the robot’s multi-stage actuation system. As reported in the table, the model-free adaptive predictive control (MFAPC) provided in [24] was also applied to the studied SSHDR. While it also did not require any modeling information and demonstrated high robustness against snowy surface conditions, the performance of the NN-RMFC, particularly in terms of steady-state error, surpassed it. Additionally, MFAPC required the implementation of multiple algorithms, including adaptive control, model predictive control, and PID control, which made its implementation somewhat challenging. On the other hand, the most common model-free control method, PID, was also applied. Although its implementation was straightforward for the studied SSHDR, it exhibited weak performance under the highly nonlinear conditions of the snowy surface. In addition, the proposed controller in [25], a fractional-order prescribed-time controller (FOPTC), was implemented in the studied SSHDR under the same conditions. The performance of the FOPTC-implemented SSHDR was competitive with the NN-RMFC, but it required a long settling time. Unfortunately, with predefinition of a better performance, FOPTC could theoretically extend indefinitely, which led to crashes in the computing system during the experiment.

### v2

@@ I Introduction
The past few years have been marked by significant global changes driven by events such as a worldwide pandemic, supply chain disruptions, resource shortages, and climate change. These factors have accelerated technological shifts, particularly through advancements in robotization and artificial intelligence (AI), which are transforming various business and industrial sectors [machadopath]. One such sector undergoing transformation is the heavy-duty machinery industry, which now requires new insights into autonomy, reliability, and process efficiency for its future development [khan2022overview]. Robotizing heavy-duty machinery requires advanced solutions that encompass both theoretical developments and application-oriented approaches, along with real-world experimental validations, to enhance automation and safety. As a result, the development of skid-steering heavy-duty robots (SSHDRs) without a separate steering mechanism has gained increasing demand in industries such as construction, agriculture, and mining. These robot platforms with differential drive systems can perform smallest radius turns by moving one set of wheels forward and the other in reverse, allowing them to pivot within their own length. This capability makes them highly maneuverable and ideal for compact, powerful, and agile operations in confined spaces, but also prone to sliding [aguilera2016general][rigotti2020mathcal]. For instance, a four-wheel SSHDR prototype operating on rough terrain is shown in Fig. 1.
@@
These SSHDRs commonly involve a continuous, real-time engagement of different stages within a framework [liikanen2019path], as shown in Fig. 2.
@@
Dynamic controllers are the lowest executive system, responsible for robustly adjusting the multi-stage mechanical actuator mechanism to generate sufficient motion and ensuring stability and goal achievement. The actuation systems integrated into SSHDRs typically consist of multi-component mechanisms designed to transform energy from the power source into mechanical motion. This process involves energy excitation, high-torque motor operation, high-ratio gearing, and, ultimately, the generation of wheel movement. The inherent complexity and nonlinearities of this multi-stage actuation system pose significant challenges in accurately modeling the SSHDR’s dynamics [shahna2024model]. Consequently, designing a model-based controller is significantly complicated by substantial modeling uncertainties and external disturbances. Hence, traditional model-free controllers remain widely used in the heavy-duty machinery industry due to their simplicity and lower computational demands, despite advancements in intelligent control methodologies. However, the performance of the common model-free controllers in higher-than-second-order systems is often limited, leading to oscillations, poor performance, or instability [zhao2017pid]. In addition, the safety and stability of SSHDRs on uneven terrain are significantly affected by wheel slippage, which arises from factors such as terrain irregularities, imbalanced load distribution, suboptimal wheel design, dynamic forces during movement, and limitations in the robot’s control mechanisms that may hinder timely adjustments [wang2008modeling]. Notably, despite the critical impact of wheel slippage, research on this issue across all types of heavy-duty robots has received limited attention over the past two decades [teji2023survey][shahna2024robust]. Various terrains with differing slip ratios must be thoroughly analyzed to ensure the stability of the robot’s autonomous controller before integrating it with mapping systems and obstacle-free trajectory generators. Hence, [liao2017performance] integrated a coordinated adaptive robust control scheme with a torque allocation technique to solve the chattering phenomenon and wheel slip compensation. However, it assumed the model of the studied skid-steering robot is accurately known. To improve the dynamic controller performance and tracking accuracy, [xiong2022path] developed a traditional controller, the cascaded proportional–integral–derivative (PID) method, while compensating for the sideslip angles. However, most studies have been built on the assumption that all four wheels of the studied robot can be controlled independently, which makes it easier to address the slipping issue and compensate for it.
@@
This paper addresses the problem of robust stabilization for a four-wheel SSHDR system, in which each two-wheel side can be controlled with a single actuator. The system is characterized by a highly nonlinear multi-component actuation model and operates on severely slippery terrains. In contrast to [liao2017performance] and [xiong2022path], the proposed control framework is entirely independent of the SSHDR model and interactive multi-stage actuation mechanisms. In addition, unlike [liikanen2019path], the robustness of the proposed control is guaranteed, adaptively ensuring operation on rough terrains with varying slip ratios. Within the proposed control framework, the functional modeling of the wheel dynamics is approximated using radial basis function neural networks (RBFNNs), where, in contrast to [yang2011adaptive] and [tong2011adaptive], the NN weights, approximation errors, and their norms are not required to be known. The proposed novel adaptive algorithm updates these values and is also responsible for compensating for disturbances caused by wheel slippage. By incorporating the online-updated values of uncertain functions and parameters into the control framework, for the first time in SSHDRs, model-free tracking of the commanded motion with global exponential stability is ensured. Global exponential stability, a strong form of stability in applications, implying uniform, asymptotic stability with exponential convergence rate, holds for all initial conditions. This work lays the foundation for advancing reliable autonomous navigation in heavy-duty machinery undergoing robotization, as shown in Fig. 2.
@@ II Differential Drive System Dynamics
where $i\in\{R,L\}$, $V_{i}$ is the linear velocity of each wheeled side, and $\bm{V}=[V_{R},V_{L}]^{\top}\in\mathbb{R}^{2}\rightarrow\mathbb{R}$. $U_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is the control input, $d_{i}(\cdot):\mathbb{R}\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ represents the uncertain multi-stage actuation model of the robot, $g_{i}\in\mathbb{R}^{+}$ is the unknown control coefficient due to the inertia of the robot, and $\Delta_{i}(t):\mathbb{R}\rightarrow\mathbb{R}$ represents external disturbances. The disturbance structure under slippage can be changed to $\Delta_{i}(t)=-{\mu_{s}}_{i}(s)F_{id}$, where $F_{id}\in\mathbb{R}\rightarrow\mathbb{R}$ represents external disturbance under no-slip conditions, and ${\mu_{s}}_{i}(s)=1+s_{i}$ denotes the functional slippage effect derived from [andreev2020trajectory]. The slip ratio of the wheel ($s_{i}\in\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$) can be defined as [jiang2022adaptive]:
@@ III Rbfnn-Based NN-RMFC
$\Phi_{ik}\left(\bm{V}\right)=\exp\left(-\frac{\left\|\bm{V}-{\varepsilon}_{ik}\right\|^{2}}{\alpha_{ik}^{2}}\right),\qquad k=1,2,\ldots,L_{i}$ (6)
where $\varepsilon_{ik}\in\mathbb{R}^{2}$ is any given two-dimensional constant vector representing the center of the receptive field, and $\alpha_{ik}\in\mathbb{R}$ is any given constant representing the width of the Gaussian functions. The notation $\|.\|$ represents the Euclidean norm. Let us propose a novel adaptive law for tuning (4) and compensating $\Delta_{i}$ in (3), as follows:
@@
Definition 1 [murray1993robotic]. Consider a nonlinear system as follows:
$\dot{x}=f(x,t)\qquad x\left(t_{0}\right)=x_{0}\qquad x\in\mathbb{R}^{n}$ (9)
@@ IV Stability Proof
where $\rho_{i}=\min\hskip 2.84544pt[\hskip 2.84544pt{g}_{i}\gamma_{i},\hskip 5.69046pt2{\kappa_{i}}\hskip 2.84544pt]$. Considering both sides ($i:R,L$), we can define an entirely quadratic function, as follows:
@@ VI-A Simulation Results in Slippery Terrains
Building upon the work of [petrovic2022analytic], an external, independent method for self-assessment is employed in this study. Specifically, the Simscape MultibodyTM simulation environment is utilized to validate the proposed control scheme and to examine the impact of the introduced slip condition on the final solution. The simulation focuses on a four-wheel SSHDR with a differential drive system (two wheels on each side), where the selection of chassis configurations is flexible, as the governing equations are not constrained to any particular design or set of physical properties. As shown in Table I, derived from [yu2020dynamic, zhu2021effects, kristalniy2022friction, canudas1999model, zeng2020experimental], during the execution, we considered randomly non-functional values of tires’ slip ratios under different environmental conditions for Eqs. (2) and (1). Although these are approximations, they are sufficient to demonstrate the control strategy’s effectiveness.
@@
Dry asphalt 0.050.400.05-0.40 0.050.200.05-0.20
Wet asphalt 0.050.800.05-0.80 0.050.500.05-0.50
Gravel 0.050.50.05-0.5 0.050.400.05-0.40
Mud 0.050.700.05-0.70 0.050.500.05-0.50
Ice 0.050.900.05-0.90 0.050.750.05-0.75
@@
The proposed control system including Eqs. (5)–(8) was applied to the simulated model, considering slip ratios provided in Table I, while the control design parameters were chosen as follows: $\kappa_{i}=1.2$, $\epsilon_{i}=0.04$, $\sigma_{i}=11.5$, $\gamma_{i}=1.6$, $L_{i}=9$, $\varepsilon_{i,k}=[2\hskip 2.84544pt{rand}-1,2\hskip 2.84544pt{rand}-1]^{\top}$, and $\alpha_{i,k}=0.13$, where randrand outputs a variable between 0 and $1$. The platform tracking errors for different operational surfaces are shown in Fig. 8, validating the efficacy of the proposed control. This figure illustrates that the control-implemented robot exhibited a higher overshoot and longer settling time on high-slip-ratio surfaces while achieving a slightly lower steady-state error after the settling time. The zoomed-in section of the figure shows that dry asphalt had the fastest settling time, followed by gravel, mud, and wet asphalt, respectively, with the icy surface having the slowest settling time.
@@ VI-B Experimental Results in Extremely Slippery and Snowy Terrains
The control design parameters were chosen as: $\kappa_{i}=1.9$, $\epsilon_{i}=0.08$, $\sigma_{i}=17.1$, $\gamma_{i}=3.6$, $L_{i}=8$, $\varepsilon_{i,k}=[2\hskip 2.84544pt{rand}-1,2\hskip 2.84544pt{rand}-1]^{\top}$, and $\alpha_{i,k}=0.15$, where randrand outputs a variable between 0 and $1$. Fig. 11 illustrates the tracking control performance by depicting the actual measured velocity of each side, commanded by the proposed controller, along with the reference velocity generated by a human-assisted joystick. Note that a difference in the sign of the velocity indicates that the robot is pivoting, as the operator demands.
@@
The comparative results in Table II indicate that, overall, the proposed NN-RMFC demonstrates better performance for the robot’s multi-stage actuation system. As reported in the table, the model-free adaptive predictive control (MFAPC) provided in [zhu2024backstepping] was also applied to the studied SSHDR. While it also did not require any modeling information and demonstrated high robustness against snowy surface conditions, the performance of the NN-RMFC, particularly in terms of steady-state error, surpassed it. Additionally, MFAPC required the implementation of multiple algorithms, including adaptive control, model predictive control, and PID control, which made its implementation somewhat challenging. On the other hand, the most common model-free control method, PID, was also applied. Although its implementation was straightforward for the studied SSHDR, it exhibited weak performance under the highly nonlinear conditions of the snowy surface. In addition, the proposed controller in [ge2023prescribed], a fractional-order prescribed-time controller (FOPTC), was implemented in the studied SSHDR under the same conditions. The performance of the FOPTC-implemented SSHDR was competitive with the NN-RMFC, but it required a long settling time. Unfortunately, with predefinition of a better performance, FOPTC could theoretically extend indefinitely, which led to crashes in the computing system during the experiment.